# Compositional Design of Safety Controllers for Large-scale Stochastic Hybrid Systems

### v1

@@ 6. Computation of CBC
Suppose sets $X_{0_{i}},X_{u_{i}},X_{i},U_{i},$ and $W_{i}$ can be described by $X_{0_{i}}=\{x_{i}\in{\mathbb{R}}^{n_{i}}\mid g_{0_{i}}(x_{i})\geq 0\}$, $X_{u_{i}}=\{x_{i}\in{\mathbb{R}}^{n_{i}}\mid g_{u_{i}}(x_{i})\geq 0\}$, $X_{i}=\{x_{i}\in{\mathbb{R}}^{n_{i}}\mid g_{i}(x_{i})\geq 0\}$, $U_{i}=\{U_{i}\subseteq{\mathbb{R}}^{m_{i}}\mid g_{\nu_{i}}(\nu_{i})\geq 0\}$, and $W_{i}=\{W_{i}\subseteq{\mathbb{R}}^{p_{i}}\mid g_{w_{i}}(w_{i})\geq 0\}$. Let there exist an SOS polynomial ¯$\bar{\mathcal{B}}_{i}(x_{i})$, constants $\gamma_{1_{i}}\in{\mathbb{R}},\bar{\kappa}_{i},\gamma_{2_{i}}\in{\mathbb{R}}^{+}$, $\bar{\mu}_{i},\bar{\beta}_{i},\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}},\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\in{\mathbb{R}}_{0}^{+}$, polynomials $l_{\nu_{j_{i}}}(x_{i})$, $\hat{l}_{\nu_{j_{i}}}(x_{i})$ corresponding to the $j^{\text{th}}$ input in $\nu_{i}=(\nu_{1_{i}},\nu_{2_{i}},\ldots,\nu_{m_{i}})\in U_{i}\subseteq{\mathbb{R}}^{m_{i}}$ and $\bar{\nu}_{i}=(\bar{\nu}_{1_{i}},\bar{\nu}_{2_{i}},\ldots,\bar{\nu}_{m_{i}})\in U_{i}\subseteq{\mathbb{R}}^{m_{i}}$, respectively. Suppose there exist vectors of SOS polynomials $l_{0_{i}}(x_{i})$, $l_{u_{i}}(x_{i})$, $\tilde{l}_{i}(x_{i})$, $l_{i}(x_{i},\nu_{i},w_{i}),\hat{l}_{i}(x_{i},\nu_{i},w_{i}),l_{\nu_{i}}(x_{i},\nu_{i},w_{i}),\hat{l}_{\nu_{i}}(x_{i},\nu_{i},w_{i}),l_{w_{i}}(x_{i},\nu_{i},\\ w_{i}),\hat{l}_{w_{i}}(x_{i},\nu_{i},w_{i})$ of appropriate dimensions such that the following expressions are SOS polynomials:
@@
¯$\bar{\mathcal{B}}_{i}(x_{i})\!-\!l_{u_{i}}^{\top}(x_{i})g_{u_{i}}(x_{i})\!-\!\bar{\beta}_{i},$ (6.1c)
-- ¯$\mathcal{L}\bar{\mathcal{B}}_{i}(x_{i})\!-\!\gamma_{1_{i}}\!(\bar{\mathcal{B}}_{i}(x_{i})\!)\!+\!\bar{\varphi}_{i_{1}}w_{i}^{\top}w_{i}\!+\!\bar{\eta}_{1_{i}}\!-\!\!\!\sum_{j=1}^{m_{i}}(\nu_{j_{i}}\!-\!l_{\nu_{j_{i}}}(x_{i})\!)-l_{i}^{\top}(x_{i},\nu_{i},w_{i})g_{i}(x_{i})$
$-l_{\nu_{i}}^{\top}(x_{i},\nu_{i},w_{i})g_{\nu_{i}}(\nu_{i})-l_{w_{i}}^{\top}(x_{i},\nu_{i},w_{i})g_{w_{i}}(w_{i}),$ (6.1d)
-- $\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(f_{2_{i}}(x_{i},\nu_{i},w_{i})\!)\big{|}x_{i},\!\nu_{i},\!w_{i}\Big{]}\!+\!\gamma_{2_{i}}\bar{\mathcal{B}}_{i}(x_{i})\!+\!\bar{\varphi}_{i_{2}}w_{i}^{\top}w_{i}+\bar{\eta}_{2_{i}}\!-\!\sum_{j=1}^{m_{i}}(\nu_{j_{i}}\!-\!\hat{l}_{\nu_{j_{i}}}(x_{i})\!)$
$-\hat{l}_{i}^{\top}(x_{i},\nu_{i},w_{i})g_{i}(x_{i})-\hat{l}_{\nu_{i}}^{\top}(x_{i},\nu_{i},w_{i})g_{\nu_{i}}(\nu_{i})-\hat{l}_{w_{i}}^{\top}(x_{i},\nu_{i},w_{i})g_{w_{i}}(w_{i}).$ (6.1e)

### v2

@@ 6 Computation of CBC
Suppose sets $X_{0_{i}},X_{u_{i}},X_{i},U_{i},$ and $W_{i}$ can be described by $X_{0_{i}}=\{x_{i}\in{\mathbb{R}}^{n_{i}}\mid g_{0_{i}}(x_{i})\geq 0\}$, $X_{u_{i}}=\{x_{i}\in{\mathbb{R}}^{n_{i}}\mid g_{u_{i}}(x_{i})\geq 0\}$, $X_{i}=\{x_{i}\in{\mathbb{R}}^{n_{i}}\mid g_{i}(x_{i})\geq 0\}$, $U_{i}=\{U_{i}\subseteq{\mathbb{R}}^{m_{i}}\mid g_{\nu_{i}}(\nu_{i})\geq 0\}$, and $W_{i}=\{W_{i}\subseteq{\mathbb{R}}^{p_{i}}\mid g_{w_{i}}(w_{i})\geq 0\}$. Let there exist an SOS polynomial $\bar{\mathcal{B}}_{i}(x_{i})$, constants $\gamma_{1_{i}}\in{\mathbb{R}},\bar{\kappa}_{i},\gamma_{2_{i}}\in{\mathbb{R}}^{+}$, $\bar{\mu}_{i},\bar{\beta}_{i},\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}},\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\in{\mathbb{R}}_{0}^{+}$, polynomial $l_{\nu_{j_{i}}}(x_{i})$, corresponding to the $j^{\text{th}}$ input in $\nu_{i}=(\nu_{1_{i}},\nu_{2_{i}},\ldots,\nu_{m_{i}})\in U_{i}\subseteq{\mathbb{R}}^{m_{i}}$. Suppose there exist vectors of SOS polynomials $l_{0_{i}}(x_{i})$, $l_{u_{i}}(x_{i})$, $\tilde{l}_{i}(x_{i})$, $l_{i}(x_{i},\nu_{i},w_{i}),\hat{l}_{i}(x_{i},\nu_{i},w_{i}),l_{\nu_{i}}(x_{i},\nu_{i},w_{i}),l_{w_{i}}(x_{i},\nu_{i},w_{i}),\\ \hat{l}_{w_{i}}(x_{i},\nu_{i},w_{i})$ of appropriate dimensions such that the following expressions are SOS polynomials:
@@
$\bar{\mathcal{B}}_{i}(x_{i})\!-\!l_{u_{i}}^{\top}(x_{i})g_{u_{i}}(x_{i})\!-\!\bar{\beta}_{i},$ (6.1c)
- $\mathcal{L}\bar{\mathcal{B}}_{i}(x_{i})\!-\!\gamma_{1_{i}}\!(\bar{\mathcal{B}}_{i}(x_{i})\!)\!+\!\bar{\varphi}_{i_{1}}w_{i}^{\top}w_{i}\!+\!\bar{\eta}_{1_{i}}\!-\!\!\!\sum_{j=1}^{m_{i}}(\nu_{j_{i}}\!-\!l_{\nu_{j_{i}}}(x_{i})\!)$
$-l_{i}^{\top}(x_{i},\nu_{i},w_{i})g_{i}(x_{i})\!-\!l_{\nu_{i}}^{\top}(x_{i},\nu_{i},w_{i})g_{\nu_{i}}(\nu_{i})$
$-l_{w_{i}}^{\top}(x_{i},\nu_{i},w_{i})g_{w_{i}}(w_{i}),$ (6.1d)
- $\mathds{E}\Big[\bar{\mathcal{B}}_{i}(f_{2_{i}}(x_{i},\nu_{i},w_{i})\!)\big|x_{i},\!\nu_{i},\!w_{i}\Big]\!+\!\gamma_{2_{i}}\bar{\mathcal{B}}_{i}(x_{i})\!+\!\bar{\varphi}_{i_{2}}w_{i}^{\top}w_{i}$
$+\bar{\eta}_{2_{i}}\!-\!\sum_{j=1}^{m_{i}}(\nu_{j_{i}}\!-\!{l_{\nu_{j_{i}}}(x_{i})}\!)-\hat{l}_{i}^{\top}(x_{i},\nu_{i},w_{i})g_{i}(x_{i})$
$-{l_{\nu_{i}}^{\top}(x_{i},\nu_{i},w_{i})}g_{\nu_{i}}(\nu_{i})-\hat{l}_{w_{i}}^{\top}(x_{i},\nu_{i},w_{i})g_{w_{i}}(w_{i}).$ (6.1e)